# MD-NOMAD: Mixture density nonlinear manifold decoder for emulating stochastic differential equations and uncertainty propagation

### v1

@@ Abstract
We propose a neural operator framework, termed mixture density nonlinear manifold decoder (MD-NOMAD), for stochastic simulators. Our approach leverages an amalgamation of the pointwise operator learning neural architecture nonlinear manifold decoder (NOMAD) with mixture density-based methods to estimate conditional probability distributions for stochastic output functions. MD-NOMAD harnesses the ability of probabilistic mixture models to estimate complex probability and the high-dimensional scalability of pointwise neural operator NOMAD. We conduct empirical assessments on a wide array of stochastic ordinary and partial differential equations and present the corresponding results, which highlight the performance of the proposed framework.
@@ 1 Introduction
Differential equations (DE) hold a pivotal position in engineering and scientific endeavors, constituting fundamental tools for inquiry and practical application. These equations serve as primary vehicles for elucidating the behavior of intricate physical systems through mathematical formulations. These formulations typically encapsulate the inherent dynamics and relationships within the systems under examination, often presenting a concise and condensed representation. Moreover, they furnish a framework essential for forecasting the intricate interactions within these systems. Even though DEs are extensively used across various scientific disciplines, obtaining analytical solutions, particularly for partial differential equations (PDEs), is often challenging and usually unattainable in practical engineering contexts marked by complex geometries, nonlinearities, and boundary conditions. Consequently, for solving PDEs in such scenarios, one resorts to numerical analysis [12, 23] with simulators based on numerical methods such as finite difference [18], finite element [25], and finite volume [17] techniques, which are employed to approximate the solutions for these equations. However, even when employing these methods, solving a PDE can evolve into a complex and time-intensive undertaking, especially in cases where complete knowledge of all parameters and boundary or initial conditions (BCs/ICs) is lacking. Furthermore, the incorporation of simplifications and assumptions to address gaps in the understanding of partially known parameters and BCs/ICs (epistemic uncertainty) inevitably might result in numerical approximations that deviate from the physical reality. Moreover, many physical systems [28, 14] inherently entail uncertainty that cannot be fully captured by deterministic models, and via extension, by deterministic simulators. However, this challenge is usually addressed by employing stochastic differential equations (SDEs), which are commonly utilized to articulate the governing physics and integrate associated uncertainty by incorporating both deterministic and stochastic influences.
While stochastic differential equations (SDEs) furnish us with the ability to model systems characterized by uncertainty, such modeling necessitates the utilization of stochastic simulators — computational models employing numerical techniques tailored for SDEs [14]. Furthermore, due to the intrinsic randomness embedded within SDEs, the output of these simulators, with a fixed set of input parameters, emerges as a random variable adhering to an unknown conditional distribution. Moreover, to describe this output distribution, stochastic simulators must be repeatedly evaluated for the same input parameters to obtain the associated statistics and probability density functions (PDFs). Consequently, while SDEs provide a framework for tackling uncertainty in physical systems, the computational expense involved in numerically modeling these equations is increased in comparison to deterministic DEs.
@@
To tackle the challenge of computational cost associated with stochastic simulators, the construction of surrogate models, which are adept at approximating the input-output relationships while maintaining computational efficiency, for these simulators has emerged as a prudent and pragmatic option. The field of surrogate modeling has garnered appreciable attention in the recent past. Within the framework of strategies inclined towards traditional methods for surrogate modeling of stochastic simulators, diverse methodologies have been investigated and refined. These include generalized additive models [11], generalized linear models [22], [7], kernel density estimators [10, 8], Gaussian processes [4], polynomial chaos expansions [2], and more sophisticated models that build upon and refine standard surrogate modeling approaches such as generalized lambda distribution-based models [35, 34] or stochastic polynomial chaos expansions [36]. However, certain limitations are associated with these methods: some are susceptible to the curse of dimensionality [32], others necessitate data to conform to specific requirements [2], some have intricate implementations [34], while others demand numerous replications for identical input parameters [35]. Additionally, certain methods impose restrictions on the shapes of response distributions that can be emulated [11, 22]. Furthermore, with the advancement of machine learning, contemporary deep learning-based methods for surrogate modeling of stochastic simulators have drawn noticeable focus. These approaches encompass various techniques, such as normalizing flow-based models [21], generative adversarial networks (GANs) [33], neural stochastic differential equations [13], surrogate modeling through optimal transport and projection pursuit [3], and gray-box methodologies like combining conditional generative moment matching networks (CGMMN) [26] with partially known SDEs [5], or integrating neural networks (NNs) with numerical integrator-motivated loss functions to identify drift and diffusivity functions in SDEs [6]. Nonetheless, it’s worth noting that CGMMN and GAN-based models can pose challenges during training, and normalizing flow-based models require significant computational resources [3].
@@
A more recent development in deep learning and surrogate modeling for DEs has been the introduction of NN architectures termed neural operators [15, 19, 29, 20, 31], which, by utilization of available data, can approximate the mapping between two continuous functions that belong to infinite-dimensional function spaces. Furthermore, the usage of some of the neural operator architectures is limited to specific types of meshes [19, 31], while others are mesh-independent and leverage a pointwise approximation strategy [29, 20]. Additionally, Fourier and Wavelet neural operators [19, 31] utilize discrete convolutions, which incur escalating computational costs as spatial dimensions increase. However, mesh-independent pointwise architectures such as NOMAD [29] and DeepONet [20] establish a separation between sensor measurements/parametric inputs and spatial/temporal/spatiotemporal location by supplying the former and latter to different sub-networks, while simultaneously linking the two subnetworks through some type of reconstruction procedure to finally obtain the output at the supplied spatial/temporal/spatiotemporal location. Moreover, the pointwise nature of NOMAD and DeepONet also aids in scalability to higher dimensional problems. However, while the aforementioned neural operator architectures, originally implemented as deterministic models, are more suitable as surrogate models for non-stochastic DEs, probabilistic variants of neural operators such as variational autoencoding NOMAD [29] and variational Bayes DeepONet [9] have also been introduced. Nevertheless, variational methodology-based deep learning models encounter numerous challenges including imprecise likelihood (unless invertible NNs are used), complexity in modeling intricate probability distributions, and training instabilities. To this end, we propose Mixture Density NOMAD (MD-NOMAD), a novel framework that blends the formalism of mixture density modeling [1] with the neural operator framework NOMAD., and functions as a surrogate model/operator learning framework for both stochastic ordinary and partial differential equations. Furthermore, the key characteristics of MD-NOMAD include:
@@
  • MD-NOMAD is easy to implement, employs standard optimization techniques, i.e, does not require specialized interventions such as reparameterization trick or stochastic gradient variational Bayes estimator, and is computationally less expensive and easy to train as compared to generative models such as normalizing flows, GANs, and variational networks.
@@
  • Since the PDFs of the employed mixture distribution are typically accessible, the requirement for employing Monte Carlo (MC) methods to estimate the statistics and PDF of the conditional output distribution during the inference stage is eliminated, and these can be computed analytically once the parameters of the mixture components are obtained from MD-NOMAD for a given input.
  • MD-NOMAD can be used to construct a surrogate model for one-shot uncertainty propagation for DEs.
The paper is structured as follows: Section 2 outlines the problem statement, while Section 3 offers an overview of the necessary background for comprehending the mathematical foundations of the proposed framework, including a brief review of mixture density networks and NOMAD. Section 4 delves into the details of the proposed framework. In Section 5, we present four examples to illustrate the performance of the proposed framework. Finally, Section 6 presents the concluding remarks.
@@ 2 Problem statement
In the context of pointwise mapping, considering a point $\bm{x}$ within the domain $\mathcal{D}$, the output function $G(f)(\bm{x})$ produces a real value, i.e., $d_{G}=1$. Additionally, in deterministic simulations, identical $\bm{x}$ and $f(\bm{x})$ inputs consistently result in the same output function values. However, with probabilistic or stochastic simulators, the uncertainty arising from the inherent characteristics of the system or insufficient knowledge thereof leads to varying output function values for a given $\bm{x}$ and $f(\bm{x})$. Therefore, to characterize the behavior of the output function for a stochastic simulator, random variable $\bm{\omega}\in\Omega$ has to be introduced. Based on this, the nonlinear differential operator for stochastic simulators can be expressed as follows:
@@
$(f(\bm{x}),\bm{\omega})$ $\mapsto G_{\bm{\omega}}(f)(\bm{x}),$ (2)
where $\Omega$ signifies the event space of the probability space $\{\Omega,\Sigma,\mathbb{P}\}$, encapsulating the randomness, and $G_{\omega}(f)(x)$ is stochastic output function. Compared to deterministic simulations, stochastic simulators incur higher computational costs as numerous simulator runs have to be performed for a given set of inputs for a good characterization of PDFs for the simulator’s output. Therefore, this paper aims to devise a mixture density-based operator learning framework for the construction of surrogate models tailored for stochastic simulators.
@@ 4 Proposed framework
Figure 1: Schematic diagram depicting the architecture of MD-NOMAD. Here, $\mathbf{x_{b}}$ serves as the input to the branch network, which is a fully connected network (FCN) producing a finite-dimensional latent feature representation as its output. The sensor values or locations $\mathbf{x_{d}}$ (also referred to as augmenting input) are augmented to the latent vector obtained from the branch network and provided as an input to the decoder network, which is also an FCN. The output of the decoder network comprises the parameters of the $\mathscr{m}$-component mixture of a probability distribution, with the probability distribution chosen from a suitable parametric family. In this article, we use a mixture of Gaussian distribution. Subsequently, the MD-NOMAD model is trained by minimizing the sum of the negative logarithm of the probability $P(G_{\omega}(\mathbf{x_{b}})(\mathbf{x_{d}}))$ over $\mathscr{n}$ training samples.
@@
As discussed in Section 3, while a deterministic NOMAD can effectively learn a mapping from input function evaluations to output function values, it cannot directly model the mapping for a stochastic operator, which typically necessitates the estimation of the PDF of the output function. Hence, to address this, we combine the pointwise operator learning NOMAD framework with the mixture density-based formalism to arrive at the mixture density NOMAD (MD-NOMAD) architecture. This MD-NOMAD framework, by design, enables the estimation of the conditional probability distribution for a stochastic output function $G_{\omega}(f)(\bm{x})$ associated with the input function $f(\bm{x})$ given a sensor value $\bm{x}$. While various parametric distributions could be considered for the mixture components, for this study, we focus solely on the Gaussian distribution due to its simplicity and interoperability. Thus, within this framework, an analog to $\rho_{i}(\bm{x};\boldsymbol{\theta})$ in Equation 3 is represented as $\rho_{\theta}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$, denoting a Gaussian distribution with mean $\mu_{\boldsymbol{\theta}}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$ and standard deviation $\sigma_{\theta}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$ (or variance $\sigma_{\boldsymbol{\theta}}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})^{2}$) for the $i$-th mixture component, where the mean, standard deviation/variance, along with the mixing coefficient, are parameterized by a NOMAD with parameters $\boldsymbol{\theta}$. In addition, to ensure that the standard deviation/variance is non-negative, softplus activation function is applied to the NOMAD predicted component parameters for standard deviation/variance. Furthermore, $\mathbf{x_{b}}$ and $\mathbf{x_{d}}$ represent the inputs to the branch and decoder network, respectively. A schematic of the MD-NOMAD’s architecture is illustrated in Figure 1.
@@ 5 Results and discussion
In addition, we showcase the utility of our model for one-shot uncertainty propagation within a setting based on 2-dimensional stochastic partial differential equations (SPDEs).

### v2

@@ Abstract
We propose a neural operator framework, termed mixture density nonlinear manifold decoder (MD-NOMAD), for emulating stochastic simulators. Our framework leverages an amalgamation of the pointwise operator learning neural architecture nonlinear manifold decoder (NOMAD) with versatile mixture density-based methods for approximating the conditional probability distributions of stochastic output functions. Furthermore, MD-NOMAD harnesses the ability of probabilistic mixture models to estimate complex probability distributions and inherits the high-dimensional scalability of NOMAD. A key feature of MD-NOMAD is its ability to perform one-shot uncertainty propagation, eliminating the need for Monte Carlo methods in statistical computations for uncertainty propagation tasks. We conduct empirical assessments on a wide array of stochastic ordinary and partial differential equations and present the corresponding results, which highlight the performance of the proposed framework.
@@ 1 Introduction
Mechanics underpins the behavior of complex physical systems and is typically modeled using differential equations (DEs). However, analytical solutions—especially for partial differential equations (PDEs)—are rarely feasible in real-world settings, leading to reliance on numerical methods such as finite difference [18], finite element [25], and finite volume [17] techniques. These methods can be computationally intensive, particularly when system parameters or boundary conditions are uncertain. Moreover, many physical systems involve inherent randomness, which deterministic models cannot capture. Stochastic differential equations (SDEs) address this by incorporating both deterministic and random effects. However, SDE modeling requires stochastic simulators — computational models employing numerical techniques tailored for SDEs. Due to randomness, for fixed inputs, the outputs of stochastic simulators are random variables with unknown conditional distributions, which necessitate repeated simulations to estimate statistics or probability density functions (PDFs). A key application of SDEs is uncertainty propagation (UP) [37], which quantifies how input uncertainties affect the model outputs. Traditional methods like Monte Carlo (MC) simulations [38], though widely used for UP problems, require a large number of samples—even with advanced variants [40]—resulting in high computational costs, especially in high-dimensional settings.
@@
To tackle the challenge of computational cost associated with stochastic simulators, the construction of surrogate models, which are adept at approximating the input-output relationships while maintaining computational efficiency, for these simulators has emerged as a prudent and pragmatic option. The field of surrogate modeling has garnered appreciable attention in the recent past. Within the framework of strategies inclined towards traditional methods for surrogate modeling of stochastic simulators, diverse methodologies have been investigated and refined. These include generalized additive models [11], generalized linear models [22], [7], kernel density estimators (KDE) [10, 8], Gaussian processes [4], polynomial chaos expansions [2], and more sophisticated models that build upon and refine standard surrogate modeling approaches such as generalized lambda distribution-based models [35, 34] or stochastic polynomial chaos expansions (SPCE) [36]. However, certain limitations are associated with these methods: some are susceptible to the curse of dimensionality [32, 36, 10, 4], some necessitate data to conform to specific requirements [2], while others demand numerous replications for identical input parameters [35]. Additionally, certain methods impose restrictions on the shapes of response distributions that can be emulated [11, 22, 34]. With the advancement of machine learning, contemporary deep learning-based methods for surrogate modeling of stochastic simulators have drawn noticeable focus. These approaches encompass various techniques, such as normalizing flow-based models [21], generative adversarial networks (GANs) [33], neural stochastic differential equations [13], denoising diffusion probabilistic models (DDPMs) [39], surrogate modeling through optimal transport and projection pursuit [3], and gray-box methodologies like combining conditional generative moment matching networks (CGMMN) [26] with partially known SDEs [5], or integrating neural networks (NNs) with numerical integrator-motivated loss functions to identify drift and diffusivity functions in SDEs [6]. Nonetheless, it’s worth noting that CGMMN and GAN-based models can pose challenges during training, and normalizing flow-based models require significant computational resources [3].
@@
A more recent development in deep learning and surrogate modeling for DEs has been the introduction of NN architectures termed neural operators [15, 19, 29, 20, 31], which, by utilization of available data, can approximate the mapping between two continuous functions that belong to infinite-dimensional function spaces. The usage of some of the neural operator architectures is limited to specific types of meshes [19, 31], while others are mesh-independent and leverage a pointwise approximation strategy [29, 20]. Fourier and Wavelet neural operators [19, 31] utilize discrete convolutions, which incur escalating computational costs as spatial dimensions increase. However, mesh-independent pointwise architectures such as NOMAD [29] and DeepONet [20] separate sensor measurements or parametric inputs from spatial, temporal, or spatiotemporal locations. They achieve this by feeding these two types of inputs into different sub-networks. These sub-networks are then linked through a reconstruction procedure to obtain the output. This process allows the model to produce outputs at the specified spatial, temporal, or spatiotemporal locations. Moreover, the pointwise nature of NOMAD and DeepONet also aids in scalability to higher dimensional problems. The aforementioned neural operator architectures were originally implemented as deterministic models. As such, they are more suitable as surrogate models for non-stochastic differential equations. To effectively handle stochastic simulators, we introduce a new framework — mixture density nonlinear manifold decoder (MD-NOMAD). This pointwise framework combines mixture density modeling [1] with the neural operator architecture NOMAD. MD-NOMAD acts as an operator learning-based surrogate model for both stochastic ordinary and partial differential equations. More importantly, although various generative models — such as SPCEs, GANs, and DDPMs — can function as surrogates for stochastic simulators, they inherently operate as sample generators during the inference stage. Consequently, MC methods remain necessary to obtain the desired statistical quantities of interest. The computation of these statistics is particularly crucial for tasks such as uncertainty propagation. In contrast, the proposed MD-NOMAD framework entirely obviates the need for MC-based computations, and therefore, enables uncertainty propagation to be performed in a one-shot manner, which helps in considerable reduction in associated computational cost. Furthermore, one of MD-NOMAD’s subnetworks — the branch network — provides a latent representation for a given parametric setting, which can be leveraged for latent space-based modeling tasks. However, the current study does not explore the utilization of the latent space. The salient features of the proposed MD-NOMAD framework are enumerated below:
@@
  • MD-NOMAD can act as a surrogate model for stochastic systems with a large number of components while employing a mixture of univariate Gaussian distributions. This is achieved by incorporating a suitable conditioning variable, which allows the decoder network of MD-NOMAD to learn the marginal distribution of individual components. This functionality of MD-NOMAD is showcased through its application to an $N$-component Lorenz-96 system.
  • Since the PDFs of the employed mixture distribution are typically accessible, the requirement for employing MC methods to estimate the statistics and PDF of the conditional output distribution during the inference stage is eliminated, and these can be computed analytically once the parameters of the mixture components are obtained from MD-NOMAD for a given input.
  • MD-NOMAD can be used to construct a surrogate model capable of performing one-shot uncertainty propagation for DEs. This capability is demonstrated on a high-dimensional elliptic stochastic partial differential equation (SPDE).
The paper is structured as follows: Section 2 outlines the problem statement, while Section 3 offers an overview of the necessary background for comprehending the mathematical foundations of the proposed framework, including a brief review of mixture density networks and NOMAD. Section 4 delves into the details of the proposed framework. In Section 5, we present six examples to illustrate the performance of the proposed framework. Finally, Section 6 presents the concluding remarks.
@@ 2 Problem statement
In the context of pointwise mapping, considering a point $\bm{x}$ within the domain $\mathcal{D}$, the output function $G(f)(\bm{x})$ produces a real value, i.e., $d_{G}=1$. Additionally, in deterministic simulations, identical $\bm{x}$ and $f(\bm{x})$ inputs consistently result in the same output function values. However, with probabilistic or stochastic simulators, the uncertainty arising from the inherent characteristics of the system or insufficient knowledge thereof leads to varying output function values for a given $\bm{x}$ and $f(\bm{x})$. Therefore, to characterize the behavior of the output function for a stochastic simulator, random variable $\boldsymbol{\omega}\in\Omega$ has to be introduced. Based on this, the nonlinear differential operator for stochastic simulators can be expressed as follows:
@@
$(f(\bm{x}),\boldsymbol{\omega})$ $\mapsto G_{\boldsymbol{\omega}}(f)(\bm{x}),$ (2)
where $\Omega$ signifies the event space of the probability space $\{\Omega,\Sigma,\mathbb{P}\}$, encapsulating the randomness, and $G_{\boldsymbol{\omega}}(f)(\bm{x})$ is stochastic output function. Compared to deterministic simulations, stochastic simulators incur higher computational costs as numerous simulator runs have to be performed for a given set of inputs for a good characterization of PDFs for the simulator’s output. Therefore, this study aims to devise a mixture density-based operator learning framework for the construction of surrogate models tailored for stochastic simulators.
Another objective of this study is to leverage the developed surrogate model to efficiently address the UP problem. Consider an input function $f_{\boldsymbol{\omega}}:\mathcal{D}\times\Omega\rightarrow\mathbb{R}^{d_{f}}$, where $\boldsymbol{\omega}\sim p(\boldsymbol{\omega})$ represents parametric uncertainty. Consequently, the stochastic output function is defined as $G(f_{\boldsymbol{\omega}})(\bm{x},\boldsymbol{\omega})$. Note that this notation diverges from that in Equation 2, as the stochasticity is introduced through the input function; however, this scenario constitutes a specific instance of the more general formulation presented in Equation 2. Within this stochastic framework, the UP problem seeks to propagate uncertainty through the nonlinear differential operator $\mathcal{W}_{s}$ and to characterize the statistical properties of the resulting output $G(f_{\boldsymbol{\omega}})(\bm{x},\boldsymbol{\omega})$. This involves tasks such as computing the expected value of the stochastic output,
[…]
determining the variability of the stochastic output about its mean,
[…]
and constructing the PDF of the output by marginalizing over the input randomness,
[…]
where $\bm{y}$ denotes a specific realization of the stochastic output function. Solving this problem using MC methods is computationally expensive due to the extensive number of simulator or surrogate runs required to obtain adequate samples. Therefore, this work proposes a one-shot strategy employing the devised mixture density-based operator learning framework to solve the UP problem. This approach obviates the need for MC methods, thereby enabling a more tractable and efficient uncertainty propagation.
@@ 4 Proposed framework
Figure 1: Schematic diagram depicting the architecture of MD-NOMAD. Here, $\mathbf{x_{b}}$ serves as the input to the branch network, which is an FCN producing a finite-dimensional latent feature representation as its output. The sensor values or locations $\mathbf{x_{d}}$ (also referred to as augmenting input) are augmented to the latent vector obtained from the branch network and provided as an input to the decoder network, which is also an FCN. The output of the decoder network comprises the parameters of the $\mathscr{m}$-component mixture of a probability distribution, with the probability distribution chosen from a suitable parametric family. In this article, we use a mixture of univariate Gaussian distribution. Subsequently, the MD-NOMAD model is trained by minimizing the sum of the negative logarithm of the probability $P(G_{\omega}(\mathbf{x_{b}})(\mathbf{x_{d}}))$ over $\mathscr{n}$ training samples.
@@
As discussed in Section 3, while a deterministic NOMAD can effectively learn a mapping from input function evaluations to output function values, it cannot directly model the mapping for a stochastic operator, which typically necessitates the estimation of the PDF of the output function. Hence, to address this, we combine the pointwise operator learning NOMAD framework with the mixture density-based formalism to arrive at the mixture density NOMAD (MD-NOMAD) architecture. This MD-NOMAD framework, by design, enables the estimation of the conditional probability distribution for a stochastic output function $G_{\omega}(f)(\bm{x})$ associated with the input function $f(\bm{x})$ given a sensor value $\bm{x}$. Although a variety of parametric distributions can be used for the mixture components, the choice largely depends on factors such as the nature of the response distribution being modeled. Some distribution families are better suited for capturing skewness or heavy tails, while others may offer advantages in terms of training stability, interpretability, or simplicity. In this study, we limit our focus to the univariate Gaussian distribution, chosen for its balance of simplicity, robustness during training, and interpretability. Thus, within this framework, an analog to $\rho_{i}(\bm{x};\boldsymbol{\theta})$ in Equation 6 is represented as $\rho_{\theta}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$, denoting a Gaussian distribution with mean $\mu_{\boldsymbol{\theta}}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$ and standard deviation $\sigma_{\theta}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})$ (or variance $\sigma_{\boldsymbol{\theta}}^{i}(\mathbf{x_{b}})(\mathbf{x_{d}})^{2}$) for the $i$-th mixture component, where the mean, standard deviation/variance, along with the mixing coefficient, are parameterized by a NOMAD with parameters $\boldsymbol{\theta}$. In addition, to ensure that the standard deviation/variance is non-negative, softplus activation function is applied to the NOMAD predicted component parameters for standard deviation/variance. Furthermore, $\mathbf{x_{b}}$ and $\mathbf{x_{d}}$ represent the inputs to the branch and decoder network, respectively. A schematic of the MD-NOMAD’s architecture is illustrated in Figure 1.
@@ 5 Results and discussion
In addition, we showcase the utility of our model for one-shot uncertainty propagation within a setting based on a 2-dimensional SPDE. The primary objective of this study is to efficiently obtain the PDF of a system’s response, thereby enabling a comprehensive characterization of its stochastic behavior and informed decision-making. In contrast to DDPM, the proposed MD-NOMAD framework offers a faster analytical approach for estimating the response PDF for any given input conditioning. Additional details regarding this comparative computational efficiency are provided in Appendix  Appendix A. A similar rationale underlies the relatively slower performance of SPCE in deriving the response density. Therefore, we benchmark the performance of MD-NOMAD against a KCDE-based model for the problems under investigation. Although KCDE, akin to SPCE, is affected by the curse of dimensionality, it possesses the advantage of not requiring sampling prior to density estimation. The conditional response distribution in KCDE is estimated as follows:
$\hat{f}_{Y|\bm{X}}(y\mid\bm{x})=\frac{\sum\limits_{i=1}^{n}K_{\bm{X}}\left(\bm{x},\bm{x}_{i},h_{x}\right)K_{Y}\left(y,y_{i},h_{y}\right)}{\sum\limits_{j=1}^{n}K_{\bm{X}}\left(\bm{x},\bm{x}_{j},h_{x}\right)},$ (12)
[…]
$\mathcal{E}_{W}=\mathbb{E}_{\bm{X}}\left[l^{W}_{2}(Y_{\bm{X}},\hat{Y}_{\bm{X}})\right],$ (13)
where $Y_{\bm{x}}$ represents the simulator response, $\hat{Y}_{\bm{x}}$ denotes the response of the surrogate model, and $l^{W}_{2}$ is the squared 2-Wasserstein distance, expressed as:
[…]
with $Q_{1}$ and $Q_{2}$ representing the quantile functions of the random variables $Y_{1}$ and $Y_{2}$, respectively. We also employ KL-divergence ($\mathcal{E}_{KL}$) [41] as an error metric. Both of these error metrics are calculated on an unseen test set for each of the example problems, utilizing both the MD-NOMAD and KCDE-based model as emulators, and the results are detailed in Table 3. Additionally, the hyperparameters associated with MD-NOMAD, including the number of mixture components, number of layers, and learning rate, are determined through a grid search for each problem.